### v1

# Beyond Freeze-Out: A Novel Freeze-in Mechanism for Dark Matter via Supercooled Phase Transitions

Abstract
We investigate a novel freeze-in mechanism for Weakly Interacting Massive Particles (WIMPs), facilitated by a supercooled first-order phase transition (FOPT) in the early universe. Unlike the conventional freeze-out and freeze-in scenarios, this mechanism allows WIMPs to acquire their relic abundance through a non-equilibrium production process following a rapid entropy injection. We explore multiple dark matter (DM) candidates, including vector, fermionic, and scalar-mediated models, and identify fermionic DM with a pseudoscalar mediator as the most viable candidate for this mechanism. The study demonstrates that FOPT dilutes preexisting DM density and simultaneously induces a rapid mass increase, preventing thermal re-equilibration and enabling DM production via freeze-in. Additionally, we analyze the gravitational wave (GW) signals associated with FOPT, identifying parameter regions detectable by future GW observatories such as LISA and UDECIGO. This framework offers a compelling alternative to traditional WIMP and Feebly Interacting Massive Particle (FIMP) scenarios, providing new avenues for DM model building and experimental verification.
1 Introduction
The nature of DM remains one of the most profound mysteries in modern physics[1, 2]. Despite overwhelming astrophysical and cosmological evidence for its existence, the particle identity of DM continues to elude discovery. The Standard Model of elementary particles(SM), despite its enormous success in describing nature in its most fundamental form, does not contain any candidates for DM. Therefore, the presence of models beyond the SM is essential to describe DM[3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Among the many theoretical paradigms proposed to explain DM, WIMPs have been extensively studied, primarily through the well-established freeze-out mechanism[59, 60, 61]. In this scenario, WIMPs achieve thermal equilibrium in the early universe and decouple as the universe expands and cools. However, traditional WIMP models face challenges from stringent experimental constraints, motivating the exploration of alternative production mechanisms.
The freeze-in mechanism offers a compelling alternative, wherein DM particles are never in thermal equilibrium but are produced gradually from the interactions of particles in the thermal bath[62]. Freeze-in has historically been associated with FIMPs, which interact so weakly that their production rate is suppressed, resulting in a relic density that matches observations.
The role of FOPT and its impact on DM phenomenology is very important[63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. Recently, however, a novel variation of the freeze-in mechanism has been proposed for WIMPs, where weak or moderate interaction strengths suffice due to the influence of a supercooled FOPT in the early universe[97](Freeze-in of WIMP). This scenario opens a new path for DM model building, enabling WIMPs to remain a viable candidate while resolving challenges associated with traditional freeze-out. It provides a robust framework for exploring physics beyond the SM, including connections to FOPT and GW signals. This new mechanism suggests WIMPs can also freeze-in under certain conditions:
  • A supercooled first-order phase transition dilutes preexisting DM density to near zero.
  • The phase transition increases the mass of DM significantly, leading to a large mass-to-temperature ratio. This prevents WIMPs from re-equilibrating with the thermal bath.
Therefore, in the study of DM production in the early universe, two main mechanisms are commonly considered: the traditional WIMP freeze-out and FIMP freeze-in. These mechanisms differ in their interaction strengths, thermal equilibrium conditions, and the role of phase transitions. Freeze-out describes the decoupling of once-equilibrated DM particles, leading to their relic abundance. Traditional freeze-in applies to feebly interacting particles that never reached equilibrium but were gradually produced over cosmic history. The WIMP freeze-in provides a new connection between FOPT and DM, opening up a third possibility for realizing DM besides WIMPs and FIMPs. It can be tested by combining the WIMP and GW experiments.
This work explores the WIMP freeze-in production of various DM candidates facilitated by an FOPT. The FOPT provides two essential effects: the injection of entropy that dilutes preexisting DM and a rapid mass increase for DM particles, preventing thermal equilibration. While the scalar WIMP candidate is a natural choice for demonstrating this mechanism[97], it can be extended to several other DM candidates.
The paper is organised as follows. In section. 2, we extend SM by several different models and DM candidates. In section. 3, we examine the conditions of the WIMP freeze-in scenario and present the numerical analyses of the models in light of supercooled FOPT scenario and study production of GWs in section 4, respectively. Finally, the conclusions present in section. 5.
2 The models
In this section, we present different models with different DM candidates, for which we will examine the conditions of the Freeze-in of WIMP in the next section. These models are presented in Table 2.
2.1 Vector DM with $U(1)$ symmetry
The Lagrangian of this model is as follows[54, 55]:
${\cal L}={\cal L}_{SM}-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+(D^{\prime}_{\mu}S)^{*}(D^{\prime\mu}S)-V(H,S),$ (2.1)
where ${\cal L}_{SM}$ is the SM Lagrangian without the Higgs potential term and the covariant derivative and field strength of $V_{\mu}$ are given as:
$D^{\prime}_{\mu}S=(\partial_{\mu}+ig_{v}V_{\mu})S,$
$V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}.$
The potential which is renormalizable and invariant under gauge and $Z_{2}$ symmetry is:
$V(H,S)=-\mu_{H}^{2}H^{\dagger}H-\mu_{S}^{2}S^{*}S+\lambda_{H}(H^{\dagger}H)^{2}+\lambda_{S}(S^{*}S)^{2}+\lambda_{SH}(S^{*}S)(H^{\dagger}H).$ (2.2)
Due to the imposed $Z_{2}$ symmetry under which only the dark gauge boson $V_{\mu}$ is odd, the vector field $V_{\mu}$ can be considered as a DM candidate. The $S$ field is the intermediate between the dark sector and the SM sector. After the electroweak symmetry breaking, the parameters of the model are given by:
$\nu_{2}=\frac{M_{V}}{g_{v}},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\sin\alpha=\frac{\nu_{1}}{\sqrt{\nu_{1}^{2}+\nu_{2}^{2}}}$
$\lambda_{H}=\frac{\cos^{2}\alpha M_{H_{1}}^{2}+\sin^{2}\alpha M_{H_{2}}^{2}}{2\nu_{1}^{2}}$
$\lambda_{S}=\frac{\sin^{2}\alpha M_{H_{1}}^{2}+\cos^{2}\alpha M_{H_{2}}^{2}}{2\nu_{2}^{2}}$
$\lambda_{SH}=\frac{\left(M_{H_{2}}^{2}-M_{H_{1}}^{2}\right)\sin\alpha\cos\alpha}{\nu_{1}\nu_{2}}$ (2.3)
Where the model includes three independent parameters $M_{V}$, $M_{H_{2}}$ and $g_{v}$.
2.2 Scale Invariant Vector DM with $U(1)$ symmetry
In this model, there is no mass term at the tree level, and electroweak symmetry breaking occurs at the one-loop level via the Gildener- Weinberg mechanism[98]. The Lagrangian of this model is as follows[56, 99]:
${\cal L}={\cal L}_{SM}+(D^{\prime}_{\mu}S)^{*}(D^{\prime\mu}S)-V(H,S)-\frac{1}{4}V_{\mu\nu}V^{\mu\nu},$ (2.4)
where ${\cal L}_{SM}$ is the SM Lagrangian without the Higgs potential term and
$D^{\prime}_{\mu}S=(\partial_{\mu}+ig_{v}V_{\mu})S,$
$V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}.$
The most general scale-invariant potential which is renormalizable and invariant under gauge and $Z_{2}$ symmetry is:
$V(H,S)=\frac{\lambda_{H}}{6}(H^{\dagger}H)^{2}+\frac{\lambda_{S}}{6}(S^{*}S)^{2}+2\lambda_{SH}(S^{*}S)(H^{\dagger}H).$ (2.5)
Similar to the previous model, the vector field $V_{\mu}$ is a candidate for DM and $S$ is the intermediate between the dark sector and the SM. After the symmetry breaking, we have the following constraints:
$\nu_{2}=\frac{M_{V}}{g_{v}},$ $sin\alpha=\frac{\nu_{1}}{\sqrt{\nu_{1}^{2}+\nu_{2}^{2}}}$
$\lambda_{H}=\frac{3M_{H_{1}}^{2}}{\nu_{1}^{2}}cos^{2}\alpha$ $\lambda_{S}=\frac{3M_{H_{1}}^{2}}{\nu_{2}^{2}}sin^{2}\alpha$
$\lambda_{SH}=-\frac{M_{H_{1}}^{2}}{2\nu_{1}\nu_{2}}sin\alpha\,cos\alpha,$ (2.6)
where $M_{V}$ is the mass of vector DM after symmetry breaking. Regarding one-loop effect, the dark Higgs mass is given by[56, 98]
$M_{H_{2}}^{2}=\frac{1}{8\pi^{2}\nu^{2}}\left(M_{H_{1}}^{4}+6M_{W}^{4}+3M_{Z}^{4}+3M_{V}^{4}-12M_{t}^{4}\right),$ (2.7)
where $M_{W,Z,t}$ are the masses of W, Z gauge bosons, and top quark, respectively. Constraints (2.6) severely restrict free parameters of the model up to two independent parameters of $M_{V}$ and $g_{v}$.
2.3 Fermionic DM with $U(1)$ symmetry
This model is also scale invariant and symmetry breaking occurs at the one-loop level. The model contains three new fields: a vector field $V_{\mu}$, a Dirac fermion field $\psi$ that can play the role of DM and a complex scalar $S$, mediates between SM and the dark sector. In the model $V_{\mu}$, $\psi$ and $S$ are charged under a new dark $U(1)_{D}$ gauge group and all of these fields are singlet under SM gauge groups. The $U(1)_{D}$ charge of the new particles are given in Table 1. The Lagrangian of this model is as follows[53]:
${\cal L}={\cal L}_{SM}+i\bar{\psi}_{L}\gamma^{\mu}D_{\mu}\psi_{L}+i\bar{\psi}_{R}\gamma^{\mu}D_{\mu}\psi_{R}-g_{s}\bar{\psi}_{L}\psi_{R}S+h.c.-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+(D_{\mu}S)^{*}(D^{\mu}S)-V(H,S).$ (2.8)
where ${\cal L}_{SM}$ is the SM Lagrangian without the Higgs potential term, The covariant derivative is
$D_{\mu}=(\partial_{\mu}+iQg_{v}V_{\mu}),\quad\text{and}$
$V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu},$ (2.9)
where $Q$ is the charge of the new particles under $U(1)$ symmetry. The most general scale-invariant potential $V(H,S)$ that is renormalizable and invariant under gauge symmetry is
$V(H,S)=\frac{1}{6}\lambda_{H}(H^{\dagger}H)^{2}+\frac{1}{6}\lambda_{S}(S^{*}S)^{2}+2\lambda_{SH}(S^{*}S)(H^{\dagger}H).$ (2.10)
After the symmetry breaking, we have the following constraints:
$\nu_{2}=\frac{M_{V}}{g_{v}},$ $sin\alpha=\frac{\nu_{1}}{\sqrt{\nu_{1}^{2}+\nu_{2}^{2}}}$
$M_{\psi}=\frac{g_{s}M_{V}}{\sqrt{2}g_{v}}$ $\lambda_{H}=\frac{3M_{H_{1}}^{2}}{\nu_{1}^{2}}cos^{2}\alpha$
$\lambda_{S}=\frac{3M_{H_{1}}^{2}}{\nu_{2}^{2}}sin^{2}\alpha$ $\lambda_{SH}=-\frac{M_{H_{1}}^{2}}{2\nu_{1}\nu_{2}}sin\alpha\,cos\alpha,$
where $M_{\psi}$ and $M_{V}$ are the masses of vector and fermion fields after symmetry breaking. Meanwhile, since we do not assume $M_{V}<2M_{\psi}$, the only candidate for DM is $\psi$. Regarding one-loop effect, the dark Higgs mass is given by[98]:
$M_{H_{2}}^{2}=\frac{1}{8\pi^{2}\nu^{2}}\left(M_{H_{1}}^{4}+6M_{W}^{4}+3M_{Z}^{4}+3M_{V}^{4}-12M_{t}^{4}-4M_{\psi}^{4}\right).$ (2.12)
where $M_{W,Z,t}$ are the masses of W, Z gauge bosons, and top quark, respectively. In the equations above, $H_{1}$ is the Higgs boson and $M_{H_{1}}=125$ GeV. According to (LABEL:2-11), the independent parameters of the model are $M_{V}$, $M_{\psi}$ and the coupling $g_{v}$
Field $S$ $\psi_{L}$ $\psi_{R}$
$U(1)_{D}$ charge 1 $\frac{1}{2}$ $-\frac{1}{2}$
Table 1: The charges of the dark sector particles under the new $U(1)_{D}$ symmetry.
2.4 Fermionic DM with scalar mediator vs. pseudoscalar mediator
In this model, the SM is extended with two new fields: a Dirac fermion $\psi$ which plays the role of the DM candidate and a (pseudo)scalar field $S$ that interacts with the DM through a Yukawa term. The Lagrangian of the model is as follows[100]:
${\cal L}={\cal L}_{SM}+{\cal L}_{dark}(\psi,S)+{\cal L}_{S}(S)+{\cal L}_{int}(H,S,\psi),$ (2.13)
where ${\cal L}_{SM}$ stands for the SM Lagrangian. The Lagrangian of fermionic DM, ${\cal L}_{dark}(\psi,S)$, is :
${\cal L}_{dark}(\psi,S)=\bar{\psi}(i\gamma_{\mu}\partial^{\mu}-M_{\psi})\psi.$ (2.14)
${\cal L}_{S}$ is as follows:
${\cal L}_{S}=\frac{1}{2}(\partial^{\mu}S)^{2}-\frac{1}{2}\mu_{S}^{2}S^{2}-\frac{1}{4}\lambda_{S}S^{4},$ (2.15)
and ${\cal L}_{int}$ is the (pseudo)scalar interaction with the dark and the SM sectors. When $S$ is a pseudoscalar then,
${\cal L}_{int}(H,S,\psi)=g_{d}S\bar{\psi}\gamma^{5}\psi+\frac{1}{2}\lambda_{SH}S^{2}H^{\dagger}H,$ (2.16)
and when $S$ represents the scalar,
${\cal L}_{int}(H,S,\psi)=g_{d}S\bar{\psi}\psi+\frac{1}{2}\lambda_{SH}S^{2}H^{\dagger}H.$ (2.17)
The Higgs potential in the SM sector reads:
$V_{H}=-\mu_{H}^{2}H^{\dagger}H-\lambda_{H}(H^{\dagger}H)^{2}.$ (2.18)
The tree-level potential which is given by the substitution $H=(0~{}~{}h\equiv\langle h_{1}\rangle)^{\dagger}$ and $S\equiv\langle h_{2}\rangle$ reads,
$V_{0}(h_{1},h_{2})=-\frac{1}{2}\mu_{H}^{2}h_{1}^{2}-\frac{1}{2}\mu_{S}^{2}h_{2}^{2}+\frac{1}{4}\lambda_{H}h_{1}^{4}+\frac{1}{4}\lambda_{S}h_{2}^{4}+\frac{1}{2}\lambda_{SH}h_{1}^{2}h_{2}^{2}.$ (2.19)
After the symmetry breaking both Higgs particle and the scalar field undergo non-zero $vev$. In fact, by substituting $h_{1}\longrightarrow\nu_{1}+h_{1}$ and $h_{2}\longrightarrow\nu_{2}+h_{2}$, the fields $h_{1}$ and $h_{2}$ mix with each other and they can be rewritten by the mass eigenstates $H_{1}$ and $H_{2}$ as
$H_{1}=h_{1}\cos\theta+h_{2}\sin\theta,$
$H_{2}=-h_{1}\sin\theta+h_{2}\cos\theta.$ (2.20)
In the above equations, $\theta$ is the mixing angle. In the following we assume that $H_{1}$ is the eigenstate of the SM Higgs with $M_{H_{1}}=125$ GeV ($\nu_{1}=246$ GeV) and $H_{2}$ corresponds to the eigenstate of the singlet scalar. After symmetry breaking, we have
$\lambda_{H}=\frac{M_{H_{2}}^{2}\sin^{2}\theta+M_{H_{1}}^{2}\cos^{2}\theta}{2\nu_{1}^{2}}$
$\lambda_{S}=\frac{M_{H_{2}}^{2}\cos^{2}\theta+M_{H_{1}}^{2}\sin^{2}\theta}{2\nu_{2}^{2}}$
$\lambda_{SH}=\frac{\left(M_{H_{2}}^{2}-M_{H_{1}}^{2}\right)\sin 2\theta}{2\nu_{1}\nu_{2}}$ (2.21)
According to the above equations, the independent parameters of the model can be chosen as $M_{H_{2}}$,$M_{\psi}$,$\theta$, $g_{d}$ and $\nu_{2}$. Note that all we have said in this section is true for both scalar and pseudoscalar mediators.
Models Symmetry Key Parameters
Vector DM with $U(1)$ symmetry $U(1)$ $M_{V},M_{H_{2}},g_{v}$
Scale Invariant Vector DM with $U(1)$ symmetry $U(1)$ $M_{V},g_{v}$
Fermionic DM with $U(1)$ symmetry $U(1)$ $M_{V},M_{\psi},g_{v}$
Fermionic DM with Scalar Mediator $Z_{2}$ $M_{H_{2}},M_{\psi},\theta,g_{d},\nu_{2}$
Fermionic DM with Pseudoscalar Mediator $Z_{2}$ $M_{H_{2}},M_{\psi},\theta,g_{d},\nu_{2}$
Table 2: Summary of DM models, their symmetries, and key parameters.
3 Freeze-in of WIMP
Freeze-in of WIMP is a novel scenario for producing DM that is the boundary between freeze-out and freeze-in mechanisms[97]. In this mechanism, in the conventional thermal history of the Universe, DM particles inevitably thermalize and freeze-out. Freeze-in of WIMPs can happen if the Universe experiences a supercooled FOPT. A supercooled FOPT releases a huge amount of entropy, which dilutes the preexisting DM density to a negligible level. The WIMPs gain mass from the FOPT and never return to equilibrium due to their large mass-to-temperature ratio. After the FOPT, DM will be accumulatively produced via the process of $SM~{}SM\rightarrow DM~{}DM$, which is a typical freeze-in scenario, but it applies to weak or moderate couplings. The main point of this introduced mechanism is that the coupling between the dark side and the SM is larger than the conventional freeze-in case.
In the early universe, DM($X$) was massless and $z\equiv(M_{X}/T)=0$. With the start of supercooling at temperature $T_{2}$, the DM mass undergoes a sudden change to $M_{X}\gg T_{2}$, where $T_{2}$ denotes the temperature after the FOPT. We assume a supercooled FOPT such that $T_{2}\gg T_{1}$. $T_{1}$ is the temperature at which the FOPT begins. This leads to an enormous increase in entropy density by a factor of $(T_{2}/T_{1})^{3}$ and dilutes the preexisting DM density to a negligible level. Consequently, the evolution of the density of DM begins at $z_{2}=M_{X}/T_{2}\gg 1$, with an initial condition $Y_{X}(z_{2})\approx(T_{1}/T_{2})^{3}Y_{1}\sim 0$. From here on, the production of DM particles begins with the freeze-in mechanism and with a negligible initial amount.
With the above interpretations, the two basic conditions for establishing the presented scenario are as follows:
  • Reheating temperature $T_{2}$ must be greater than the critical temperature $T_{1}$($T_{2}\gg T_{1}$):
    In this scenario, $T_{2}$(reheating temperature) must always be greater than $T_{1}$. $T_{\Lambda}$ is the temperature at which supercooling(starting a phase of thermal inflation) begins and when reheating is prompt ($\Gamma\gg H$), we have[101, 63]:
    $T_{2}=T_{\Lambda}=(\frac{30V_{\Lambda}}{\pi^{2}g_{*}})^{\frac{1}{4}}.$ (3.1)
    In the above equation, $g_{*}\approx 106.75$ and for all models presented in the paper, $\Gamma\gg H$, where $\Gamma$ is the decay width of scalars and $H$ is the Hubble constant during the FOPT. The $V_{\Lambda}$ quantity is also defined as follows[63]:
    $V_{\Lambda}=\frac{\beta_{\lambda_{S}}v_{2}^{4}}{16},$ (3.2)
    where $v_{2}$ is the vacuum expectation value of the scalar mediator $S$ and $\beta_{\lambda_{S}}$ is the one-loop beta function for coupling of scalar mediator $\lambda_{S}$. In the following, we examine this condition for all the presented models.
    1. 1.
      The Model 2.1 :
      This medel includes three independent parameters $M_{V}$, $M_{H_{2}}$ and $g_{v}$. For this model, all possible phase transition states have been investigated in reference [102]. In this case, there is no phase transition, either one-step or two-step, that can be observed in consistent with the DM constraints, including the correct relic density of DM[103]. Therefore, this model cannot be included in the mechanism proposed for the production of DM in this paper.
      Refer to caption     Refer to caption
                 (a)                                                                     (b)
      Refer to caption     Refer to caption
             (c)                                                                     (d)
      Figure 1: Temperature variations for the model 2.2 with respect to DM mass for different couplings.
      Refer to caption     Refer to caption
                 (a)                                                                     (b)
      Figure 2: Temperature variations for the model 2.3 with respect to DM mass. In (a) $M_{V}=500$ GeV and in (b) $g_{v}=1$ is considered. We know that in the model must always $M_{\psi}<M_{V}/2$ so that the only DM particle in the model is the fermion $\psi$.
      Refer to caption     Refer to caption
                 (a)                                                                     (b)
      Figure 3: Temperature variations with respect to the independent parameters of the model 2.4. The mixing angle is assumed $\theta=1^{\circ}$. In (a) $v_{2}=200$ GeV, in (b) $M_{H_{2}}=5500$ GeV and in (c) $M_{H_{2}}=5500$ GeV and $v_{2}=200$ GeV is considered.
    2. 2.
      The Model 2.2 :
      This medel includes two independent parameters $M_{V}$ and $g_{v}$. According to equation 2.7, $M_{V}$ must always be greater than 240 GeV. Figure 1 shows the behavior of critical temperature($T_{1}$) and reheating temperature($T_{2}$) for different couplings and masses. As can be seen, only for a very small portion of the parameter space(240 GeV$<M_{V}<$ 500 GeV), $T_{2}$ is larger than $T_{1}$. But to establish the mechanism of Freeze-in of WIMP, we must assume that $T_{2}\gg T_{1}$, which is not the case here, and both temperatures are very close and $T_{2}/T_{1}\sim 1$. Therefore, this model cannot be included in the framework of the presented mechanism.
    3. 3.
      The Model 2.3 :
      This medel includes three independent parameters $M_{V}$, $M_{\psi}$ and $g_{v}$. First, it is important to note that in the model, for our only candidate for DM to be a fermionic particle($\psi$), we must have $M_{\psi}<M_{V}/2$. In Figure 2(a), temperature changes are plotted against the mass of DM for constant $g_{v}$ and $M_{V}$. As can be seen, $T_{2}\sim T_{1}$ and $T_{2}/T_{1}\sim 1$. But from $M_{\psi}>250$ GeV, $T_{2}$ will be a relatively large distance from $T_{1}$, which is not the case in the model scenario where only the fermionic particle is a candidate for DM($M_{\psi}<M_{V}/2=250$ GeV). In Figure 2(b), temperature changes are plotted for different $M_{V}$. As can be seen, as $M_{V}$ increases, $T_{2}$ becomes smaller than $T_{1}$. But for smaller $M_{V}$, $T_{2}$ becomes a relatively significant distance from $T_{1}$ at a point where the single-component DM condition(only $\psi$ as a DM candidate) in the model is violated(because of $M_{\psi}>M_{V}/2$). Therefore, this model like the previous two models cannot be included in the framework of the presented mechanism.
    4. 4.
      The Model 2.4 :
      Since the model 2.4 with scalar mediator is very limited and it has been publicly ruled out by direct discovery experiments[100], we examine Model 2.4 with pseudoscalar mediator. By choosing the mediator in the model 2.4 to be a pseudoscalar[100], the elastic scattering cross section for DM-nucleon is velocity suppressed and so the theory easily evades the direct detection bounds from XENONnT and LZ[104, 105]. There are five independent parameters $M_{H_{2}}$,$M_{\psi}$,$\theta$, $g_{d}$ and $\nu_{2}$ in this model. An overall result of the Higgs signal strength measured by ATLAS and CMS will restrict the Higgs mixing angle to values smaller than $\sin\theta\leq 0.12$ $(\theta\leq 6.9^{\circ})$[106]. Therefore, in the rest of the paper we consider the mixing angle$(\theta=1^{\circ})$. Also, the mass of DM($M_{\psi}$) will not have any effect on determining the temperatures $T_{1}$ and $T_{2}$. In Figure 3(a), temperature changes are plotted against the mass of $H_{2}$ for constant $g_{d}$ and $\nu_{2}$. As can be seen, as the $M_{H_{2}}$ decreases, $T_{2}$ will become much larger than $T_{1}$, and $\frac{T_{2}}{T_{1}}\gg 1$. In Figure 3(b), temperature changes are plotted against the $\nu_{2}$ for constant $g_{d}$ and $M_{H_{2}}$. As can be seen, as the $\nu_{2}$ decreases, $T_{2}$ will become much larger than $T_{1}$. In Figure 3(c), temperature changes are plotted against the $g_{d}$ for constant $\nu_{2}$ and $M_{H_{2}}$. As can be seen, $T_{2}$ is always greater than $T_{1}$, but their changes are constant with $g_{d}$. As a result, the two determining parameters for examining temperatures are $M_{H_{2}}$ and $\nu_{2}$. Therefore, this model has the initial condition $T_{2}\gg T_{1}$ for the mechanism of freeze-in of WIMP and allows for the examination of this mechanism. It can also be protected from strong constraints of direct detection.
      In Figure 4, the allowed parameter space is also plotted in consistent with the constraint of $T_{2}>T_{1}$ for independent parameters of model. A summary of the analysis performed in this section is provided in Table 3.
So according the above points, we conclude that only fermionic dark matter with a pseudo scalar mediator satisfies the condition of $T_{2}\gg T_{1}$. The supercooled FOPT enables freeze-in WIMP production specifically for it due to a combination of unique conditions and constraints that are satisfied only in this scenario. Here’s a detailed explanation:
  1. 1.
    Entropy Injection and Dilution of Preexisting DM: A supercooled FOPT releases a significant amount of entropy into the universe. This entropy injection dilutes any preexisting DM density to nearly zero. This is a crucial condition for the freeze-in mechanism, as it ensures that the DM production starts from a negligible initial abundance. For fermionic DM with a pseudoscalar mediator, this dilution is particularly effective because the pseudoscalar mediator allows for a specific type of interaction that can maintain the DM out of thermal equilibrium after the phase transition.
  2. 2.
    Mass Increase and Suppression of Thermal Equilibration: During the FOPT, the mass of the DM particles increases significantly. This large mass-to-temperature ratio prevents the DM particles from re-equilibrating with the thermal bath after the phase transition. In the case of fermionic DM with a pseudoscalar mediator, the mass increase is such that the DM particles remain out of equilibrium, which is essential for the freeze-in mechanism to operate. The pseudoscalar mediator facilitates this by providing a coupling that is weak enough to prevent re-equilibration but strong enough to allow for gradual DM production.
Models Condition of $T_{2}\gg T_{1}$
Vector DM with $U(1)$ symmetry Impossible
Scale Invariant Vector DM with $U(1)$ symmetry Impossible
Fermionic DM with $U(1)$ symmetry Impossible
Fermionic DM with Scalar Mediator Impossible
Fermionic DM with Pseudoscalar Mediator Possible
Table 3: Condition of $T_{2}\gg T_{1}$ for different models
  • Failure to return to equilibrium conditions after FOPT ($T_{2}<T_{dec}$):
    In this scenario, for $T_{2}>T_{dec}$ , the plasma thermalizes again, and the usual freeze-out mechanism yields the relic abundance[101, 63]. $T_{dec}\sim M_{DM}/25$ is temperature of freeze-out. So we are looking for points in the parameter space where $T_{2}<T_{dec}$. Consequently, the evolution of Boltzmann equation begins at $z_{2}=M_{DM}/T_{2}\gg 1$ where the equation is as follows:
    $\frac{dY_{X}}{dz}=-\sqrt{\frac{\pi g^{2}_{*,s}}{45g_{*}}}\frac{M_{Pl}M_{X}}{z^{2}}\langle\sigma\nu_{rel}\rangle(Y_{X}^{2}-Y_{eq}^{2}),$ (3.3)
    where $Y_{X}=n_{X}/s$ is the yield of $X$ with $s$ being the entropy density, $z=M_{X}/T$, $M_{Pl}=1.22\times 10^{19}$ GeV is the Planck scale, $g_{*,s}$ and $g_{*}$ are the numbers the relativistic degrees of freedom for entropy and energy, respectively. $\sigma\nu_{rel}$ is the thermal average of the annihilation cross section multiplying the relative velocity $\nu_{rel}$ under the Maxwell-Boltzmann distribution, and $K_{i}(z)$ is the $i$-th modified Bessel function. Finally, the allowable parameter space according to the Freeze-in WIMP mechanism is plotted in Figure 5. These points are in consistent with the observed relic density of DM[103]. The conditions of $T_{2}\gg T_{1}$ and $T_{2}<T_{dec}$ hold for all points. For these points, $z_{2}$ is between 26 and 32. Of course, we know that $M_{\psi}$ is the mass of DM.
Figure 4: Allowed parameter space in agreement with the condition of $T_{2}>T_{1}$ for the model 2.4 with pseudoscalar mediator . The mixing angle is assumed $\theta=1^{\circ}$. Our scan is done with 4000 GeV$<M_{H_{2}}<$ 6000 GeV, 100 GeV$<\nu_{2}<$ 500 GeV, and the dark coupling in the range 0.01 $<g_{d}<$ 2.
Figure 5: The parameter space that realizes the WIMP freeze-in scenario for model with pseudoscalar mediator. These points are in consistent with the observed relic density of DM, $T_{2}\gg T_{1}$ and $T_{2}<T_{dec}$. The mixing angle is assumed $\theta=1^{\circ}$. Our scan is done with 7 TeV$<M_{\psi}<$ 10 TeV, 100 GeV$<\nu_{2}<$ 200 GeV, 4500 GeV$<M_{H_{2}}<$ 6000 GeV and the dark coupling in the range 0.01 $<g_{d}<$ 2.
4 Gravitational wave signals
The total thermal effective potential is as follows[53]:
$V_{eff}=V_{0}(h_{1},h_{2})+V_{1-loop}(h_{1},h_{2};0)+V_{1-loop}(h_{1},h_{2};T),$ (4.1)
where $V_{0}$ is the tree-level potential in eq.2.19, $V_{1-loop}(h_{1},h_{2};0)$ is the Coleman-Weinberg one-loop correction at zero temperature [107] and $V_{1-loop}(h_{1},h_{2};T)$ is the one-loop thermal correction[108]. In the high-temperature approximation, the one-loop effective potential takes the following form:
$V_{1-loop}^{high-T}(h_{1},h_{2};T)=(\frac{1}{2}\kappa_{h_{1}}h_{1}^{2}+\frac{1}{2}\kappa_{h_{2}}h_{2}^{2})T^{2},$ (4.2)
in which we have
$\kappa_{h_{1}}=\frac{1}{48}(9g_{1}^{2}+3g_{2}^{2}+12g_{t}^{2}+24\lambda_{H}+4\lambda_{SH}),$ (4.3)
$\kappa_{h_{2}}=\frac{1}{12}(4\lambda_{SH}+3\lambda_{S}-g_{d}^{2}).$ (4.4)
We have dropped the Colman-Wienberg zero-temperature correction since at high temperature approximation only the thermal corrections are dominant. In eq.4.3 the parameters $g_{1}$ and $g_{2}$ are respectively the $SU(2)_{L}$ and $U(1)_{Y}$ SM couplings and $g_{t}$ is the top quark coupling.
We use the potential 4.1 to calculate GWs from the model. In the case of significant supercooling, the GW signal arises from bubble collisions[101]. To calculate the GW spectra, we need the phase transition strength $\alpha$, inverse phase transition duration $\beta/H_{*}$, $v_{\omega}$ and $T_{*}(T_{2})$. The bubble collision contribution to the spectrum of GW in the our scenario is given by[109, 110]:
$\Omega_{coll}(f)h^{2}=1.67\times 10^{-5}\Bigl{(}\frac{T_{2}}{T_{\Lambda}}\Bigl{)}^{4/3}\Bigl{(}\frac{\beta}{H_{\ast}}\Bigl{)}^{-2}\Bigl{(}\frac{\kappa\alpha}{1+\alpha}\Bigl{)}^{2}\Bigl{(}\frac{g_{\ast}}{100}\Bigl{)}^{-\frac{1}{3}}\Bigl{(}\frac{0.11v_{\omega}^{3}}{0.42+v_{\omega}^{2}}\Bigl{)}S_{coll},$ (4.5)
where $S_{coll}$ parametrises the spectral shape and is given by
$S_{coll}=\frac{3.8(f/f_{coll})^{2.8}}{2.8(f/f_{coll})^{3.8}+1},$ (4.6)
where
$f_{coll}=1.65\times 10^{-5}\Bigl{(}\frac{T_{2}}{T_{\Lambda}}\Bigl{)}^{1/3}\Bigl{(}\frac{0.62}{\upsilon_{\omega}^{2}-0.1\upsilon_{\omega}+1.8}\Bigl{)}\Bigl{(}\frac{\beta}{H_{\ast}}\Bigl{)}\Bigl{(}\frac{T_{\Lambda}}{100}\Bigl{)}\Bigl{(}\frac{g_{\ast}}{100}\Bigl{)}^{1/6}Hz.$ (4.7)
The $\alpha$ and $\beta/H_{*}$ parameters are as follows[111]:
$\alpha=\frac{\Delta\Bigl{(}V_{eff}-T\frac{\partial V_{eff}}{\partial T}\Bigl{)}\bigg{|}_{T_{*}}}{\rho_{\ast}},$ (4.8)
$\frac{\beta}{H_{\ast}}=T_{*}\frac{d}{dT}\Bigl{(}\frac{S_{3}(T)}{T}\Bigl{)}\bigg{|}_{T_{*}},$ (4.9)
where $\rho_{\ast}$ is
$\rho_{\ast}=\frac{\pi^{2}g_{\ast}}{30}T_{N}^{4}.$ (4.10)
The velocity of the bubble wall($v_{\omega}$) which is anticipated to be close to 1 for the strong transitions[112] and $g_{*}\sim 100$ is the effective number of relativistic degrees of freedom in the thermal plasma. Figure 6 shows the gravitational wave spectrum for the parameters given in table 4. As can be seen, the peak of this spectrum will fall within the range of the LISA and UDECIGO experiments[111, 113].
$M_{\psi}$ (GeV) $g_{d}$ $\nu_{2}$ (GeV) $M_{H_{2}}$ (GeV) $T_{2}$ (GeV) $T_{1}$ (GeV) $\alpha$ $\beta/H_{\ast}$
1 8573 1.26 164.57 5198.14 281.42 1.49 $4.05\times 10^{8}$ 10
2 8390 0.34 140.96 5297.67 286.09 1.5 $2.27\times 10^{9}$ 100
Table 4: Input and output parameters for two benchmark points.
Figure 6: GW spectrum for benchmark points of the Table 4.
5 Conclusion
In the paper, we explored a generalized freeze-in mechanism for DM production, leveraging the effects of a supercooled FOPT. We demonstrated that the entropy injection and mass increase associated with FOPT create conditions that allow WIMPs to undergo freeze-in rather than the conventional freeze-out mechanism. Our study considered multiple DM candidates, including vector, fermionic, and scalar-mediated DM models, identifying scenarios where freeze-in is a viable production mechanism.
Among the models examined, we found that a fermionic DM scenario with a pseudoscalar mediator best satisfies the required conditions for WIMP freeze-in, particularly in maintaining out-of-equilibrium dynamics post-FOPT. This model provides a unique framework where DM relic abundance can be generated with moderate couplings, making it distinct from traditional freeze-in models that typically involve extremely feeble interactions.
Additionally, we analyzed the GW signals associated with the supercooled FOPT, identifying parameter regions where the predicted GW spectra fall within the sensitivity ranges of future experiments such as LISA and UDECIGO. These findings highlight the potential for upcoming GW observations to probe new physics related to DM production.
Overall, this study establishes WIMP freeze-in as a robust and unifying framework for DM model building, offering new perspectives on DM candidates beyond conventional thermal histories. Future research may focus on refining parameter constraints through collider and GW observations, further solidifying the role of FOPT in shaping the early universe’s DM dynamics.